.A bead slides on a smooth rod which is rotating about one end in a vertical plane with uniform angular velocity Omega .form the equation of motion ?


1
Expert's answer
2021-04-26T18:19:20-0400

Let the mass of the bead be m and it is rotating with angular velocity "\omega" .



Applying the conservation of energy,

"\Rightarrow L= \frac{m v^2}{2}+\frac{mr\omega^2}{2}-mgr\sin\theta"

Now, applying Lagrangian equation,

"\frac{\delta }{\delta t}(\frac{\delta L}{\delta v})-\frac{\delta L}{\delta r}=0"


"\Rightarrow \frac{\delta L}{\delta r}=mr (\dfrac{\delta\omega}{\delta t})^2 -mg\sin\theta"


"\Rightarrow \frac{\delta L}{\delta v}= m\frac{dr}{dt}"


"\Rightarrow \frac{\delta }{\delta t}(\frac{\delta L}{\delta v})=m\frac{d^2r}{dt^2}"


"m\dot{\dot{r}}=mr\omega^2-mg\sin\omega t"